OpenStudy (babyslapmafro):

Please help me with unit tangent and normal vectors. I am asked to find T(t) and N(t) at the given point t=1. (click to see equation)

OpenStudy (babyslapmafro):

\[r(t)=(t^2-1)i+tj\]

OpenStudy (babyslapmafro):

\[r'(t)=2ti+j\]

OpenStudy (babyslapmafro):

\[T(1)=\frac{ 2 }{ \sqrt5 }i+\frac{ 1 }{ \sqrt5 }j\]

OpenStudy (babyslapmafro):

I tried to find N(t) a t=1 but i am not getting the right answer

OpenStudy (babyslapmafro):

\[N(t)=\frac{ T'(t) }{ ||T'(t)|| }\]

OpenStudy (babyslapmafro):

\[T(t)=\frac{ 2 }{ \sqrt5 }ti+\frac{ 1 }{ \sqrt5 }j\]

OpenStudy (babyslapmafro):

\[T'(t)=\frac{ 2 }{ \sqrt5 }\]

zepdrix (zepdrix):

\[\large \hat T(1)=\frac{<2,1>}{\sqrt5} \\ \large \hat T(1)=\left<\frac{2}{\sqrt5},\frac{1}{\sqrt5}\right>\] Bahhh I was doing N wrong, lemme try this again.. Your Tangent vector looks good though.

OpenStudy (babyslapmafro):

Ya i checked my t-vector in the back of the book but my N is way off for some reason....

zepdrix (zepdrix):

I'm a little confused by your T(t). Shouldn't we get something like this? Remember we're not plugging in t=1 at this point. \[\large \hat T(t)=\frac{<2t,1>}{\sqrt{4t^2+1}} \qquad = \qquad \left<\frac{2t}{\sqrt{4t^2+1}},\;\frac{1}{\sqrt{4t^2+1}}\right>\] And then we'd have to differentiate this ... I guess :P

OpenStudy (babyslapmafro):

Oh ok, I see my mistake now....i wasn't taking t into account when finding the magnitude of r'(t)

OpenStudy (babyslapmafro):

i think i closed this too soon lol how do you take the derivative of T(t)?

zepdrix (zepdrix):

Oh boy, it's a doozy. Quotient rule on the \(\large \hat i\) component. And for the \(\large \hat j\) component we can just use the power rule by changing the square root to a fractional exponent.

zepdrix (zepdrix):

Do you happen to have an answer key handy? I worked it out, I wanna know if I'm on the right track here.

OpenStudy (babyslapmafro):

the answer to N(1)

zepdrix (zepdrix):

ya

OpenStudy (babyslapmafro):

\[<\frac{ 1 }{ \sqrt5 },-\frac{ 2 }{ \sqrt5 }>\]

zepdrix (zepdrix):

Ah ok, I messed up somewhere then lol. I'll have to try again.

zepdrix (zepdrix):

Yesssss finally got it!

zepdrix (zepdrix):

Getting stuck at any particular spot?

OpenStudy (babyslapmafro):

ya im stuck simplifying the derivative

zepdrix (zepdrix):

That's where I was making a mistake also. Don't both simplifying it. If you've taken the derivative, simply plug in t=1 from that point.

zepdrix (zepdrix):

It will be much easier to simplify down after you've plugged in t=1.

OpenStudy (babyslapmafro):

ok and what about when im finding the magnitude of T'(t)? it doesnt need to be simplified?

zepdrix (zepdrix):

Once you have found, \(\large \hat T'(t)\), plug in t=1. From there you can simplify \(\large \hat T'(1)\) down to 2 nice looking components. Then finding \(\large ||\hat T'(1)||\) from there will be fairly straight forward.

zepdrix (zepdrix):

You'll want to use the simplified \(\large \hat T'(1)\) to find your \(\large ||\hat T'(1)||\). Don't try to find the magnitude until you've simplified the thing down! :O

zepdrix (zepdrix):

Hope that makes sense :3

OpenStudy (babyslapmafro):

ya ill just try and work it out.. this is pretty insane because this is the first problem out of 15 and I imagine each one is a bit more difficult

zepdrix (zepdrix):

Are they all the same type of problem? Yah the vector problems can take a lot of steps to get through :\

OpenStudy (babyslapmafro):

i have 12 problems like this and more on binormal vectors

zepdrix (zepdrix):

Just to help lead you in the right direction, Once you plug in t=1, and simplify your derivative, you should get this, \[\large \hat T'(1)=\left<\frac{2}{\sqrt{125}},\;-\frac{4}{\sqrt{125}}\right>\]

OpenStudy (babyslapmafro):

ok thanks